OpenStudy (anonymous):

Find the amplitude and the period of b(x) = -5sin(Πx).

OpenStudy (anonymous):

How do I solve this ?

OpenStudy (perl):

we use formula period = 2pi / |B|

OpenStudy (anonymous):

that is one ugly pi

OpenStudy (perl):

for general sinusoidal y = A sin (Bx-C) + D

OpenStudy (perl):

looks ok to me

OpenStudy (anonymous):

\[-5\sin(\pi x)\]

OpenStudy (anonymous):

\[-5\sin(\Pi x)\] either way i guess

OpenStudy (perl):

oh, its a capital Pi

OpenStudy (anonymous):

yeah

OpenStudy (perl):

im not familiar with the latex stuff but i notice cant use alt 233 , and other alt commands

OpenStudy (perl):

@mashlyashela period = 2pi / pi , in your case. so period = 2 . do you agree?

OpenStudy (anonymous):

Oh okay. So amplitude is 5 and the period is 2 ? @perl